40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 8     Entry time: Thu Nov 8 18:54:41 2007
Author: Stefan 
Type: Laser 
Category: PSL 
Subject: 35W laser ISS loop closed with LIGO ISS electronics 
I closed the ISS loop using a iLIGO type ISS board, with an additional 2kHz pole at the last stage.

I haven't carefully measured noise or OLG yet, but it is not famous yet:
Looking at the in-loop diode I get a RIN of ~3e-7/rtHz at 100Hz, and ~1e-8 at 10kHz.

The gain is limited by two sharp resonances that rise at ~40kHz when the gain is increased.

TBD:
- Fix the ISS test point wiring - these cables have 800Ohm resistance!
(No not impedance... shield to core measures 800Ohms at DC...)
This brought back painful mamories from ~3years ago when we installed
the H1 ISS... apparently the known problem wasn't fixed on all versions.
- Measure OLG, in particular check whether weird AOM transfer function still
exists - it could be killing our gain.
- Measure sensing noise in-situ
- Install sensitive OL PD
ELOG V3.1.3-